please someone help. i give feedback
Write an equation for the quadratic graphed below x-intercepts: (-1,0) and (3,0). y-intercept: (0,-3) Preview
Answer #1

1 Takes zero at (-1,0) and (3,0). . Ya acrt 1) (2-3) now at x = 0, ya -3 – 3 = a (1) (-3) .aal. ya (x+1)(x-3)